OpenStudy (anonymous):

Use the Distance Formula and x-axis of the coordinate plane. Show why the distance between two points on a number line (the x-axis) is |a-b|, where a and b are the x-coordinates of the points. So far I have:

OpenStudy (hamde):

because the distance is alength and there is no negative length..

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

\(\large d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \) \(\large d = \sqrt{(a - b)^2 + (0 - 0)^2} \) \(\large d = \sqrt{(a - b)^2 } \) \(\large d = |a - b| \)